A machine purchased three years ago for $300,000 has a current book value using straight-line depreciation of $177,000; its operating expenses are $31,000 per year. A replacement machine would cost $238,000, have a useful life of eleven years, and would require $13,000 per year in operating expenses. It has an expected salvage value of $67,000 after eleven years. The current disposal value of the old machine is $80,000; if it is kept 11 more years, its residual value would be $18,000.

Required
Calculate the total costs in keeping the old machine and purchase a new machine. Should the old machine be replaced?
Answer #1
Old machine new machine
Opportunity cost/Purchase value (80,000-18,000) = 62,000 (238,000-67,000) = 171,000
Operating cost (31,000*11) = 341,000 (13,000*11) = 143,000
Total Cost 403,000 314,000

2. Yes (as cost is higher)